A graphing iron is thrown vertically to catch on a ledge 7.5 metres above the thrower. It`s height is give in meters and time in seconds. After being thrown the equation is given by h= -4.9t?2+11t+1.5. Will the graphing iron reach the ledge?
Can you please help me? Thank you so much in advance:)

Answer by ankor@dixie-net.com(22740)

A graphing iron is thrown vertically to catch on a ledge 7.5 metres above the thrower.
It`s height is give in meters and time in seconds.
After being thrown the equation is given by h= -4.9t?2+11t+1.5.
Will the graphing iron reach the ledge?
:
A Graphing Iron? Perhaps you mean a grappling iron, anyway this is a quadratic equation, find the axis of symmetry, then the max height
x = -b/(2a) is the formula for the axis of symmetry
t = -11/(2*-4.9)
t = -11/-9.8
t = 1.12245 sec to reach max height
:
Find Max height
h = -4.9(1.12245^2) + 11(1.12245) + 1.5
h = -6.17347 + 12.347 + 1.5
h ~ 7.67 meters is max height, it should just make it.